OpenStudy (anonymous):

For breakfast, the cafeteria offers scrambled eggs with one slice of toast or an egg sandwich made with two slices of toast. In fifteen minutes, they sold 13 breakfasts and used 21 slices of bread. Use a model to determine how many egg sandwiches they sold.

OpenStudy (anonymous):

okay, so you can write a system of equations for this do you know where to start?

OpenStudy (anonymous):

not really, I have trouble with writing system of equations.

OpenStudy (anonymous):

ok, sure well let's call the number of scrambled egg meals sold a and the number sandwiches sold b, ok? then we can say that a + b = 13 because 13 meals were sold and because each sandwich uses two slices of toast, we can say that a + 2b = 21 does that make sense?

OpenStudy (anonymous):

yes, but I am still kind of confused about how to solve it. do i replace the variables with 0? Because I tried that and I don't get the right answer.

OpenStudy (anonymous):

alright, now we want to multiply one of the equations by a number that will allow us to eliminate a variable. Here, because there are the same coefficients for a, we're going to eliminate a. multiply one of the equations by -1, all the way across

OpenStudy (anonymous):

I did -1+2b=21 and got b=11 then I tried -1+b=13 and got b= 14. I also did a+-1=13 and got a = 14 and a+2(-1)=21 and got a=23. But I don't think either of those are the answer.

OpenStudy (anonymous):

no, i mean do this -1(a+b) = -1(13) now simplify

OpenStudy (anonymous):

I'm sorry for sounding stupid but i just got a+b=13 when I simplified.

OpenStudy (anonymous):

-a-b=-13 now you want to add them together like this  add up and down